OpenStudy (gabylovesyou):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x + y = 2 correct ? @ganeshie8

ganeshie8 (ganeshie8):

correct !

OpenStudy (gabylovesyou):

I am? I was totally not sure with this one lol

ganeshie8 (ganeshie8):

finally one q you're not sure ha... kets work this a bit

OpenStudy (gabylovesyou):

ok lol

ganeshie8 (ganeshie8):

(2, −2) and (−5, 12) m = \(\frac{12--2}{-5-2} = \frac{14}{-7} = -2\)

OpenStudy (gabylovesyou):

yes that is what i got for slope ..

ganeshie8 (ganeshie8):

equation of lien with slope -2, and going thru (2, -2) : \(y--2 = -2(x-2)\) \(y + 2 = -2x + 4)\) \(y = -2x + 2)\) \(y + 2x = 2)\)

ganeshie8 (ganeshie8):

you're perfectly correct !

OpenStudy (gabylovesyou):

ohhh !! thank you :)

ganeshie8 (ganeshie8):

np :)